Can anyone help me?
An experimental rocket is launched into the air from ground level with the Earth with an initial velocity of 192 feet per second. Its height s in feet is given by the formula
s=-16t^2+ 192t
After how many seconds will the rocket reach a height of 576 feet?

Answer by ewatrrr(24785)

576 = -16t^2 + 192t
16t^2 - 192t + 576 = 0
t^2 - 12t + 36 = 0
(t-6)(t-6)= 0 , t = 6sec